OpenStudy (anonymous):

find r so that the line through (2,6) and (-4,r) is perpendicular to x+2y=1

OpenStudy (tkhunny):

x + 2y = 1 ==> Slope = -1/2 (Solve for y to find Slope-Intercept Form) (r-6)/(-4-2) = 2 -- Where did the "2" come from?

OpenStudy (anonymous):

that's part of the equation how it was written in the book i am using

OpenStudy (tkhunny):

?? Did you solve it? My question was the 2 that I wrote. Where did I get get it. It was not in your book.

OpenStudy (anonymous):

no, i don't

OpenStudy (tkhunny):

Except for horizontal (Slope is zero) and vertical (Slope doesn't exist) the slopes of perpendicular lines have this relationship... \( Slope_{1}\cdot Slope_{2} = -1\) For you, the given line has slope -1/2. Substitute that for \(Slope_{1}\) and see where the algebra leads for \(Slope_{2}\)